Tohban Report 2018-02-21
From RHESSI Wiki
|Start Date:||14 Feb 2018|
|End Date:||21 Feb 2018|
|List all reports|
Very quiet week. A few B-class flares.
How many GOES flares occurred?
Flares above B, C, M, X class were 11 0 0 0
RHESSI flare list contains 0 over the time range 13-Feb-18 20-Feb-18
No overlapping RHESSI flares to count, sorry
Thanks to additional passes, SSR is down. Currently at 53.7%, with 4 passes still to go today.
Unfortunately, BGS is down today, and possibly for longer -- we'll be getting only three passes a day for some time, and the SSR is expected to creep up fast. It has been decided to turn off the very noisy D6.
Detectors 3, 6 and 8 are active and taking data.
No apparent data gaps until Feb 18th, where there are some orbit-long holes. This, and the lack of data after the 19th, is likely due to the SSR still being quite full.
The State-Of-Health packet counter flag in the RHESSI status page turned red quite often this week... Jim thinks this happens when AGO passes are not recorded properly. No adverse effects that we can observe, though.
No State-of-Health plots (those controlled by Ops) since Jan 22, 2018.
|HLAT Decimation||Rear decimation|
|Night time data (fronts)||plus/minus 4 minutes from daylight|
|Night time data (rears)||plus/minus 4 minutes from daylight|
|Require extra passes?||No|
|Requirement for moving pointer?||No|
|Attenuator operation||logic enabled|
|Detector problems?||See notes above.|